OpenStudy (anonymous):

Without using a calculator, evaluate (cos(30))^2−(sin(30))^2= 30's are degrees.

OpenStudy (raden):

use the identity : cos^2 (x) - sin^2 (x) = cos(2x)

OpenStudy (raden):

here,given x=30 deg

OpenStudy (anonymous):

oh and the 2's are supposed to be like ^2. not just multiplying.

OpenStudy (raden):

yup, that is a squared sign

OpenStudy (anonymous):

yes i know... i was just letting you know. so what you said is still right, correct?

OpenStudy (raden):

if u put x=30 to the identity above : cos^2 (x) - sin^2 (x) = cos(2x), giving us cos^2 (30) - sin^2 (30) = cos(2 * 30) = cos60, right ?

OpenStudy (anonymous):

i dont get where you are getting cos(2x)...

OpenStudy (raden):

that is one of identity in trigono, it must be in ur book, right

OpenStudy (anonymous):

uhm, i am sure it is. but we didnt learn it that way. anyway, continue.

OpenStudy (raden):

well, the alternative way did u memorize the value of sine, cos, tangen for the specials angles in trigo ?

OpenStudy (raden):

cos30 and sin30 is very familiar, because it is the special angle

OpenStudy (raden):

whatt's cos30 = ... sin30 = ...

OpenStudy (anonymous):

i have the unit circle? if that is what you mean...

OpenStudy (anonymous):

sqrt3/2 correct?

OpenStudy (anonymous):

and 1/2

OpenStudy (raden):

yup, for cos30

OpenStudy (raden):

correct again :)

OpenStudy (anonymous):

yayyyyy (:

OpenStudy (anonymous):

so sqrt3/2-1/2...

OpenStudy (raden):

now, the original ques is cos30 squared - sin30 squared = (sqrt(3)/2)^2 - (1/2)^2 = 3/4 - 1/4 = ... ?

OpenStudy (anonymous):

1/2 :D

OpenStudy (anonymous):

you are my new friend!!! thank you sooooo mush :DDD

OpenStudy (raden):

that's right it's same like cos60, what i said before :)

OpenStudy (raden):

you're welcome

OpenStudy (anonymous):

wow thats an easier way of thinking about it!

OpenStudy (raden):

actually, YES!! hhe

OpenStudy (anonymous):

this woman has us going around the world when we can just walk across the street -_-